OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I'll fan and give a medal!!

OpenStudy (abbycross167):

Find the inverse of f(x) = x3. Is the inverse a function? Explain

OpenStudy (abbycross167):

@Abbles could you please try to help me?

TheSmartOne (thesmartone):

first change f(x) to y then switch x to y, and y to an x finally solve for y :P

OpenStudy (abbycross167):

Ok I did that but I couldn't figure out how to solve for y... by the way it's f(x) = x^3 I typed it wrong in the first post

TheSmartOne (thesmartone):

Yeah, I realized. We just have to take the cube root to solve for y :P

OpenStudy (abbycross167):

so you mean, y=^3? thats the answer?

TheSmartOne (thesmartone):

for example: x = y^2 take the square root on both sides \(y = \sqrt{x}\)

jimthompson5910 (jim_thompson5910):

Following @TheSmartOne 's instructions, you'd have these steps (or similar) \[\Large f(x) = x^3\] \[\Large y = x^3\] \[\Large x = y^3\] \[\Large y^3 = x\] \[\Large y = \sqrt[3]{x}\] \[\Large f^{-1}(x) = \sqrt[3]{x}\]

OpenStudy (abbycross167):

how did you get the f^-1?

TheSmartOne (thesmartone):

f^(-1) means the inverse function of f(x)

OpenStudy (abbycross167):

oh okay, so that is the answer

jimthompson5910 (jim_thompson5910):

after you swap x and y, then solve for y, the result of solving for y leads you to the inverse of f(x). The notation f^(-1) is notation for the inverse of f(x)

OpenStudy (abbycross167):

oh alright, I think i understand. thank y'all for all the help!!

TheSmartOne (thesmartone):

Anytime! :)